OpenStudy (anonymous):

Use the Intermediate Value Theorem to verify that f(x) has a zero in the given interval. Then use the method of bisections to find an interval of length 1/32 that contains the zero.

OpenStudy (anonymous):

f(x) = x^2 - 7, [2,3]

OpenStudy (anonymous):

can u help with this last one then i'm done for today @zepdrix

zepdrix (zepdrix):

Well we can determine where f(x) has a zero by setting the function equal to zero and solving for x.\[\large f(x)=x^2-7 \qquad \rightarrow \qquad 0=x^2-7 \qquad \rightarrow \qquad x=\sqrt 7\] \(\sqrt 7\) is approximately \(2.65\) which is between 2 and 3. I'm not exactly sure how to apply the Intermediate Value Theorem, or Bisections for that matter :( Hmmm But that's where the root is located at least :C

OpenStudy (anonymous):

okay thanks for the help i appreciate it